# AdaDetectGPT: Adaptive Detection of LLM-Generated Text with Statistical Guarantees

### v1

@@ 5 Conclusion
We propose AdaDetectGPT, an adaptive LLM detector that learns a witness function $w$ to boost the performance of existing logits-based detectors. A natural approach to learning $w$ is to maximize the TNR of the resulting detector for a fixed FNR level $\alpha$. However, the TNR is a highly complicated function of $\alpha$ and $w$, which makes the learned witness function inherently $\alpha$-dependent — that is, it maximizes the TNR at a particular FNR level but does not guarantee optimality at other FNR levels. To address this, we derive a lower bound on the TNR and propose to learn $w$ by maximizing this lower bound. Notice that the lower bound separates the effects of $\alpha$ and $w$: the witness function $w$ affects the lower bound only through $T_{w}^{(2)*}$, which is independent of $\alpha$. Consequently, the witness function that maximizes this lower bound simultaneously maximizes it across all FNR levels.
In our implementation, we opted to learn the witness function via a B-spline basis due to the straightforward nature of the optimization (finding the optimal witness function boils down to solving a system of linear equations) and its favorable theoretical properties (the estimation error can be shown to attain Stone’s optimal convergence rate, Stone, 1982). The number of -spline basis functions can be selected in a data driven way for instance via Lepski’s method; see Lepski & Spokoiny (1997).
@@ 8. Answer: [Yes]
  9. Justification: We discuss the limitation of the work in Appendix H.
@@ 18. Answer: [Yes]
  19. Justification: The detailed setting for reproducibility are provided in Section 4 and Appendix F.
@@ 28. Answer: [Yes]
  29. Justification: The details for experiments are shown in Appendix F.
@@ 38. Answer: [Yes]
  39. Justification: We provided the compute resources in Appendix FHardware details.
@@ 48. Answer: [Yes]
  49. Justification: The broader impacts of this work is discussed in Appendix H.
@@ 58. Answer: [Yes]
  59. Justification: In Section 4 and Appendix F, we have explicitly cited or credited the assets used in the paper and explicitly mentioned the corresponding licenses.
@@ 60. Guidelines:
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
@@ Appendix A Additional related works: Watermarking-based detection
Our work is related to a line of research that frames watermark detection as a statistical hypothesis testing problem (see e.g., Kirchenbauer et al., 2023; Hu et al., 2024; Kuditipudi et al., 2024; Zhao et al., 2024; Li et al., 2024, 2025; Chen et al., 2025b). Under this framework, rejection of the null hypothesis (that no watermark is present) provides statistical evidence that the text was likely generated by an LLM.
@@ Appendix B Details on the analytic example in Section 3
Appendix C Implementation details
C.1 Data for estimating witness function
In this part, we illustrate how we fetch external human and machined-generate text datasets for training witness function in our experiments.
When testing the performance of AdaDetectGPT on one dataset (e.g., XSum), we recruit the remaining datasets (e.g., SQuAD and WritingPrompt) for training the witness function. This ensures the data for testing would not be included for training. In the white-box setting, since $q^{(s)}$ is known, then the machine-generated text in the training datasets are generated by $q^{(s)}$. In the black-box setting, since $q^{(s)}$ is unknown, the machine-generated text in the training datasets are generated by the surrogate models. In the case where the sampling model and scoring model is different and both using surrogate models, then the generated texts comes from the surrogate scoring model.
C.2 Details: the witness function estimation
In this section we elaborate the estimation of witness function. Recall that the population version of objective function for estimating $w$ is
$\frac{\sum_{t}[\mathbb{E}_{X_{<t}\sim p,\widetilde{X}_{t}\sim q_{t}}w(\log q_{t}(\widetilde{X}_{t}|X_{<t}))]-\sum_{t}[\mathbb{E}_{X_{<t}\sim p,\widetilde{X}_{t}\sim p_{t}}w(\log q_{t}(\widetilde{X}_{t}|X_{<t}))]}{\sqrt{\sum_{t}\mathbb{E}_{X_{<t}\sim p}\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}}(w(\log q_{t}(\widetilde{X}_{t}|X_{<t})))}},$
@@ C.2 Details: the witness function estimation
and we replace the expectations $\mathbb{E}_{X_{<t}\sim p}$ in both the numerator and denominator with their empirical average over the passages of human. This leads to the following plug-in estimated objective function:
$\frac{\sum_{i}\sum_{t}\mathbb{E}[w(\log q_{t}(\widetilde{X}_{t}|X^{(i)}_{<t}))]-\sum_{i}\sum_{t}[w(\log q_{t}(X^{(i)}_{t}|X^{(i)}_{<t}))]}{\sqrt{\sum_{i}\sum_{t}\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}}(w(\log q_{t}(\widetilde{X}_{t}|X^{(i)}_{<t})))}}$.
$\approx$ $\frac{\sqrt{2}\sum_{i}\sum_{t}\mathbb{E}[w(\log q_{t}(\widetilde{X}_{t}|X^{(i)}_{<t}))]-\sqrt{2}\sum_{i}\sum_{t}[w(\log q_{t}(X^{(i)}_{t}|X^{(i)}_{<t}))]}{\sqrt{\sum_{i}\sum_{t}\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}}(w(\log q_{t}(\widetilde{X}_{t}|X^{(i)}_{<t})))+\sum_{i}\sum_{t}\textrm{Var}(w(\log q_{t}(X^{(i)}_{t}|X^{(i)}_{<t})))}}$,
where $\widetilde{X}_{t}$ is sampled from the machine and the approximation holds because of Assumption 4. Besides, taking the expectation over $\widetilde{X}_{t}$ is time consuming, then we approximate $\mathbb{E}[w(\log q_{t}(\widetilde{X}_{t}|X^{(i)}_{<t}))$] and $\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}}(w(\log q_{t}(\widetilde{X}_{t}|X^{(i)}_{<t})))$ with
$\mathbb{E}[w(\log q_{t}(\widetilde{X}_{t}|\widetilde{X}_{<t}))]\;\textup{ and }\;\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}}(w(\log q_{t}(\widetilde{X}_{t}|\widetilde{X}_{<t})))$
where $\widetilde{X}_{<t}$ is sampled from LLM. They can be easily estimated by the Monte Carol method as we can utilize the accessible machine to generate texts. Specifically, we generate a text $\bm{\widetilde{X}}^{(i)}\sim q$. And they are estimated by:
$\mathbb{E}[w(\log q_{t}(\widetilde{X}^{(i)}_{t}|\widetilde{X}^{(i)}_{<t}))]\;\textup{ and }\;\textrm{Var}_{\widetilde{X}^{(i)}_{t}\sim q_{t}}(w(\log q_{t}(\widetilde{X}^{(i)}_{t}|\widetilde{X}^{(i)}_{<t})))$.
After these approximation, we lead to the following two-sample $t$-test type objective that eliminates the constant terms:
$\frac{\sum_{i}\sum_{t}[w(\log q_{t}(\widetilde{X}^{(i)}_{t}|\widetilde{X}^{(i)}_{<t}))]-\sum_{i}\sum_{t}[w(\log q_{t}(X^{(i)}_{t}|X^{(i)}_{<t}))]}{\sqrt{\sum_{i}\sum_{t}\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}}(w(\log q_{t}(\widetilde{X}_{t}|\widetilde{X}^{(i)}_{<t})))+\sum_{i}\sum_{t}\textrm{Var}(w(\log q_{t}(X^{(i)}_{t}|X^{(i)}_{<t})))}}$
$=$ $\frac{\sum_{i}\sum_{t}[w(\log q_{t}(\widetilde{X}^{(i)}_{t}|\widetilde{X}^{(i)}_{<t}))]-\sum_{i}\sum_{t}[w(\log q_{t}(X^{(i)}_{t}|X^{(i)}_{<t}))]}{\sqrt{\sum_{i}\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}}(\sum_{t}w(\log q_{t}(\widetilde{X}_{t}|\widetilde{X}^{(i)}_{<t})))+\sum_{i}\textrm{Var}(\sum_{t}w(\log q_{t}(X^{(i)}_{t}|X^{(i)}_{<t})))}}$
@@
For ease notations, we denoted $\log q(X^{(i)}_{t}|X^{(i)}_{<t})$ as $z^{(h)}_{it}$, which is the logit of the $t$-th token of the $i$-th human text. Similarly, for the logit computed from machine-generated text, we define it as $z^{(m)}_{it}$. Using these notations, we get the objective function below:
$\frac{1}{\sqrt{\textup{Var}(\sum_{t}w(z^{(h)}_{t}))+\textup{Var}(\sum_{t}w(z^{(m)}_{t}))}}\left(\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}w(z^{(h)}_{it})-\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}w(z^{(m)}_{it})\right)$. (10)
For simplicity, (10) implicitly assumes (i) dataset includes $n$ passages, and (ii) each passage has $L$ tokens. Actually, the computational procedure below can be easily extended to various sample sizes and various tokens number cases.
@@
Recall that the witness function have the form $w(z)=\phi(z)^{\top}\beta$, where $\phi(z)$ are some B-spline basis (De Boor, 1978) and $\beta$ be the parameter. Plug in the linear function to (10), the numerator becomes:
$\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(h)})^{\top}\beta-\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(m)})^{\top}\beta$
@@
As for the denominator, we estimate $\textup{Var}(\sum_{t}w(z^{(h)}_{t}))$ by $\beta^{\top}\widehat{\Sigma}^{(h)}\beta$, and estimate $\textup{Var}(\sum_{t}w(z^{(m)}_{t}))$ by $\beta^{\top}\widehat{\Sigma}^{(m)}\beta$, where in particular $\widehat{\Sigma}^{(h)}=\sum_{i=1}^{n}\widehat{\Sigma}_{i}^{(h)}$ with
@@
and $\widehat{\Sigma}^{(m)}=\sum_{i=1}^{n}\widehat{\Sigma}_{i}^{(m)}$; the expression for the $\widehat{\Sigma}_{i}^{(m)}$’s is analogous to that of the $\widehat{\Sigma}_{i}^{(h)}$’s and is therefore omitted. Consequently, the objective function can be rewritten as:
$\frac{\left[\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(h)})^{\top}-\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(m)})^{\top}\right]\beta}{\sqrt{\beta^{\top}(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})\beta}}$
$=$ $\left[\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(h)})^{\top}-\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(m)})^{\top}\right]\beta\times\frac{1}{\|(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})^{1/2}\beta\|_{2}}$
$=$ $\left[\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(h)})^{\top}-\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(m)})^{\top}\right]\times(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})^{-\frac{1}{2}}\alpha,$
where $\alpha=(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})^{1/2}\beta/\|(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})^{1/2}\beta\|_{2}$ and thus $\|\alpha\|_{2}$. The maximizer of the objective function has a closed-form solution:
@@
with
$\widetilde{\alpha}=(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})^{-\frac{1}{2}}\left[\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(h)})-\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(m)})\right];$ (11)
and thus $\|\widehat{\alpha}\|_{2}=1$. For $\alpha$ satisfying $\|\alpha\|_{2}=1$, we have $\beta=(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})^{-1/2}\alpha$. Therefore, the expression of $\widehat{\beta}$ is:
$\widehat{\beta}$ $=(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})^{-1/2}\widehat{\alpha}$. (12)
@@
From Section G.3, considering the increase of n_base would raise higher computational cost, and thus, we recommend to set n_base=16 and order=2.
Appendix D Assumptions for theories
In this section, we list the assumptions required for the theorems presented in Section 3 to hold, and discuss when they can reasonably be expected to hold in practice and how they may be relaxed.
D.1 Assumptions
We work under the following assumptions on the data generating process.

### v2

@@ 5 Conclusion
We propose AdaDetectGPT, an adaptive LLM detector that learns a witness function $w$ to boost the performance of existing logits-based detectors. A natural approach to learning $w$ is to maximize the TNR of the resulting detector for a fixed FNR level $\alpha$. Our proposal has two novelties. First, by connecting Fast-DetectGPT’s statistic to martingale theory and applying the MCLT, we obtain a closed-form expression for the classification threshold that achieves FNR control at level $\alpha$. Second, the TNR is a highly complex function of $\alpha$ and $w$, which makes the learned witness function $\alpha$-dependent — that is, it maximizes the TNR at a particular FNR level but does not guarantee optimality at other FNR levels. To address this, we derive a lower bound on the TNR and propose to learn $w$ by maximizing this lower bound. Our lower bound separates the effects of $\alpha$ and $w$: the witness function $w$ affects the lower bound only through $T_{w}^{(2)*}$, which is independent of $\alpha$. Consequently, the witness function that maximizes this lower bound simultaneously maximizes it across all FNR levels.
In our implementation, we opted to learn the witness function via a B-spline basis due to the straightforward nature of the optimization (finding the optimal witness function boils down to solving a system of linear equations) and its favorable theoretical properties (the estimation error can attain Stone’s optimal convergence rate, Stone, 1982, see Appendix E.6 for more details). The number of basis functions can be selected in a data driven way via Lepski’s method (Lepski & Spokoiny, 1997).
@@ 8. Answer: [Yes]
  9. Justification: We discuss the limitation of the work in Appendix G.
@@ 18. Answer: [Yes]
  19. Justification: The detailed setting for reproducibility are provided in Section 4 and Appendix C.
@@ 28. Answer: [Yes]
  29. Justification: The details for experiments are shown in Appendix C.
@@ 38. Answer: [Yes]
  39. Justification: We provided the compute resources in Appendix CHardware details.
@@ 48. Answer: [Yes]
  49. Justification: The broader impacts of this work is discussed in Appendix G.
@@ 58. Answer: [Yes]
  59. Justification: In Section 4 and Appendix C, we have explicitly cited or credited the assets used in the paper and explicitly mentioned the corresponding licenses.
@@ 60. Guidelines:
    • The name of the license (e.g., $C$-BY 4.0) should be included for each asset.
@@ Appendix A Additional related works: Watermarking-based detection
Our work is related to a line of research that frames watermark detection as a statistical hypothesis testing problem (see e.g., Kirchenbauer et al., 2023; Hu et al., 2024; Kuditipudi et al., 2024; Zhao et al., 2024; Li et al., 2025a, b; Chen et al., 2025b). Under this framework, rejection of the null hypothesis (that no watermark is present) provides statistical evidence that the text was likely generated by an LLM.
@@ Appendix B Details on the analytic example in Section 3
Appendix C Experiment details
Details for witness function estimation. In this part, we illustrate how we fetch external text datasets for training witness function in our experiments. As mentioned in the main text, when testing the performance of AdaDetectGPT on one dataset (e.g., XSum), we randomly select two other datasets (e.g., SQuAD and WritingPrompt) for training the witness function. This ensures the data for testing would not be included for training.
Recall that the population version of objective function for estimating $w$ is
$\frac{\sum_{t}[\mathbb{E}_{X_{<t}\sim p,\widetilde{X}_{t}\sim q_{t}}w(\log q_{t}(\widetilde{X}_{t}|X_{<t}))]-\sum_{t}[\mathbb{E}_{X_{<t}\sim p,\widetilde{X}_{t}\sim p_{t}}w(\log q_{t}(\widetilde{X}_{t}|X_{<t}))]}{\sqrt{\sum_{t}\mathbb{E}_{X_{<t}\sim p}\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}}(w(\log q_{t}(\widetilde{X}_{t}|X_{<t})))}}.$ (10)
In the implementation, we made three modifications to this objective function to facilitate the computation, accommodate black-box settings with unavailable logits and handle prompts that are not explicitly included in the text:
  1. 1.
@@ Appendix C Experiment details
    We replace the expectation $\mathbb{E}_{X_{<t}\sim p}$ in the numerator of (10) with its empirical average over the human-authored passages $\{\bm{X}^{(i)}\}_{i}$. This leads to the following objective function:
    $\frac{\sum_{i}\sum_{t}\mathbb{E}_{\widetilde{X}_{t}\sim q_{t}}[w(\log q_{t}(\widetilde{X}_{t}|X^{(i)}_{<t}))]-\sum_{i}\sum_{t}[w(\log q_{t}(X^{(i)}_{t}|X^{(i)}_{<t}))]}{n\sqrt{\sum_{t}\mathbb{E}_{X_{<t}\sim p}\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}}(w(\log q_{t}(\widetilde{X}_{t}|X_{<t})))}}.$ (11)
  2. 2.
    Taking the expectation $\mathbb{E}_{\widetilde{X}_{t}\sim q_{t}}$ and the variance $\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}}$ is time-consuming, as these operations need to be repeated $L$ times – once at every token position $t$. To address this, we approximate $X_{<t}^{(i)}$ in the first term of the numerator of (11) by $\widetilde{X}_{t}^{(i)}$, sampled from the LLM distribution $q_{t}$. This, in turn, allows us to approximate the expectation by
    $\frac{\sum_{i}\sum_{t}[w(\log q_{t}(\widetilde{X}_{t}^{(i)}|\widetilde{X}^{(i)}_{<t}))]-\sum_{i}\sum_{t}[w(\log q_{t}(X^{(i)}_{t}|X^{(i)}_{<t}))]}{n\sqrt{\sum_{t}\mathbb{E}_{X_{<t}\sim p}\textrm{Var}_{\widetilde{X}_{t}\sim q_{t}}(w(\log q_{t}(\widetilde{X}_{t}|X_{<t})))}},$
    where $\bm{\widetilde{X}}^{(i)}$ denotes the $i$th passage generated from $q$ by prompting the LLM to rewrite $\bm{X}^{(i)}$. As for the variance operator, we similarly approximate $\mathbb{E}_{X_{<t}\sim p}$ in the denominator by $\mathbb{E}_{X_{<t}\sim q}$. This allows us to upper bound the denominator by $n\sqrt{\sum_{t}\textrm{Var}(w(\log q_{t}(\widetilde{X}_{t}|\widetilde{X}_{<t})))}$, leading to the following lower bound of the objective function,
    $\frac{\sum_{i}\sum_{t}[w(\log q_{t}(\widetilde{X}_{t}^{(i)}|\widetilde{X}^{(i)}_{<t}))]-\sum_{i}\sum_{t}[w(\log q_{t}(X^{(i)}_{t}|X^{(i)}_{<t}))]}{n\sqrt{\sum_{t}\textrm{Var}(w(\log q_{t}(\widetilde{X}_{t}|\widetilde{X}_{<t})))}},$
    where the variance in the denominator can be estimated by the sampling variance estimator
    $\widehat{\textrm{Var}}(w(\log q_{t}(\widetilde{X}_{t}|\widetilde{X}_{<t})))=\frac{1}{n-1}\sum_{i=1}^{n}\Big[w(\log q_{t}(\widetilde{X}_{t}^{(i)}|\widetilde{X}_{<t}^{(i)}))-\frac{1}{n}\sum_{j=1}^{n}w(\log q_{t}(\widetilde{X}_{t}^{(j)}|\widetilde{X}_{<t}^{(j)}))\Big]^{2}.$
  3. 3.
    To further simplify the objective function, we interchange the order of $\sum_{t}$ and $\widehat{\textrm{Var}}$ in the denominator, leading to
    $\frac{\sum_{i}\sum_{t}[w(\log q_{t}(\widetilde{X}_{t}^{(i)}|\widetilde{X}^{(i)}_{<t}))]-\sum_{i}\sum_{t}[w(\log q_{t}(X^{(i)}_{t}|X^{(i)}_{<t}))]}{n\sqrt{\widehat{\textrm{Var}}(\sum_{t}w(\log q_{t}(\widetilde{X}_{t}|\widetilde{X}_{<t})))}}.$ (12)
    Additionally, since the numerator incorporates both human- and LLM-authored text, we refine the denominator of Equation (12) by replacing the sampling variance estimator with a simple average of the estimators computed using human- and LLM-written text
    $\frac{1}{2}\widehat{\textrm{Var}}\left(\sum_{t}w(\log q_{t}(\widetilde{X}_{t}|\widetilde{X}_{<t}))\right)+\frac{1}{2}\widehat{\textrm{Var}}\left(\sum_{t}w(\log q_{t}(X_{t}|X_{<t}))\right).$
    This yields our final objective function, in the form of a two-sample $t$-test statistic,
    $\frac{\sum_{i}\sum_{t}[w(\log q_{t}(\widetilde{X}_{t}^{(i)}|\widetilde{X}^{(i)}_{<t}))]-\sum_{i}\sum_{t}[w(\log q_{t}(X^{(i)}_{t}|X^{(i)}_{<t}))]}{n\sqrt{0.5\widehat{\textrm{Var}}(\sum_{t}w(\log q_{t}(\widetilde{X}_{t}|\widetilde{X}_{<t})))+0.5\widehat{\textrm{Var}}(\sum_{t}w(\log q_{t}(X_{t}|X_{<t})))}}.$ (13)
Compared to its population-level version (10), (13) is more suitable for black-box settings. In such settings, the distribution $q$ is unknown, making the expectation and variance over $q$ in (10) infeasible to compute. In contrast, (13) relies on text generated by the LLM. Therefore, even without access to $q$, we can still prompt the target LLM to produce $\widetilde{\bm{X}}$. Likewise, for detecting text generated under specific prompts, we can incorporate these prompts into the rewriting process to produce $\widetilde{\bm{X}}$.
@@
We next discuss the computation of $\widehat{w}$ that maximizes (13). To ease notation, we denoted $\log q(X^{(i)}_{t}|X^{(i)}_{<t})$ as $z^{(h)}_{it}$. Similarly, for the log-probabilities computed from machine-generated text, we define them as $z^{(m)}_{it}$s. Using these notations, (13) can be represented by
$\frac{1}{\sqrt{\textup{Var}(\sum_{t}w(z^{(m)}_{t}))+\textup{Var}(\sum_{t}w(z^{(h)}_{t}))}}\left(\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}w(z^{(m)}_{it})-\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}w(z^{(h)}_{it})\right)$, (14)
up to some proportional constant.
@@
Recall that we restrict the witness function to take a linear form of $w(z)=\phi(z)^{\top}\beta$, where $\phi(z)$ denotes the B-spline basis function (De Boor, 1978) and $\beta$ denotes the regression coefficients. Then numerator of (14) then becomes
$\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(m)})^{\top}\beta-\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(h)})^{\top}\beta,$
@@
whereas the denominator becomes $\sqrt{\beta^{\top}(\widehat{\Sigma}^{(m)}+\widehat{\Sigma}^{(h)})\beta}$, where $\widehat{\Sigma}^{(h)}=\sum_{i=1}^{n}\widehat{\Sigma}_{i}^{(h)}$,
@@
and $\widehat{\Sigma}^{(m)}$ can be similarly defined. Consequently, the objective function can be rewritten as:
$\frac{\left[\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(m)})^{\top}-\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(h)})^{\top}\right]\beta}{\sqrt{\beta^{\top}(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})\beta}}$
$=$ $\left[\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(m)})^{\top}-\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(h)})^{\top}\right]\beta\times\frac{1}{\|(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})^{1/2}\beta\|_{2}}$
$=$ $\left[\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(m)})^{\top}-\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(h)})^{\top}\right]\times(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})^{-\frac{1}{2}}\alpha,$
where $\alpha=(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})^{1/2}\beta/\|(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})^{1/2}\beta\|_{2}$ whose $\ell_{2}$ norm equals 1. It is immediate to see that the argmax $\widehat{\alpha}$ has a closed-form expression,
@@
where
$\widetilde{\alpha}=(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})^{-\frac{1}{2}}\left[\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(m)})-\sum_{i=1}^{n}\frac{1}{L}\sum_{t=1}^{L}\phi(z_{it}^{(h)})\right].$ (15)
This leads to
$\widehat{\beta}=(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})^{-1/2}\widehat{\alpha}$,
@@
Name Model Scale (Billion)
GPT-2 (Radford et al., 2019) openai-community/gpt2-xl 1.5B
GPT-Neo (Black et al., 2021) EleutherAI/gpt-neo-2.7B 2.7B
OPT-2.7 (Zhang et al., 2022) facebook/opt-2.7b 2.7B
GPT-J (Wang & Komatsuzaki, 2021) EleutherAI/gpt-j-6B 6B
Qwen2.5 (Yang et al., 2024a) Qwen/Qwen2.5-7B 7B
Mistral (Jiang et al., 2023) mistralai/Mistral-7B-v0.3 7B
Llama (AIMeta, 2024) meta-llama/Meta-Llama-3-8B 8B
GPT-NeoX (Black et al., 2022) EleutherAI/gpt-neox-20b 20B
{'role': 'system', 'content': 'You are a Review writer on Yelp.'},
and
respectively.
Setting on experiments with adversarial attacks. In Table 3, we have conducted experiments to evaluate the robustness of AdaDetectGPT against 2 adversarial attacks: (i) paraphrasing, where an LLM is instructed to rephrase human-written text, and (ii) decoherence, where the coherence LLM-generated text is intentionally reduced to avoid detection. These experiments were carried out across 3 datasets and 3 types of sampling and scoring models setup, resulting in a total of 18 settings. Both adversarial attacks were implemented following Bao et al. (2024).
Appendix D Technical assumptions
In this section, we list the assumptions required for the theorems presented in Section 3 to hold, and discuss when they are expected to hold and how they may be relaxed.